OpenStudy (anonymous):

a_n= (-1)^n (n+4/n+1), determine whether the sequence converges or diverges

OpenStudy (anonymous):

divergent

OpenStudy (anonymous):

how did u get that lol?

OpenStudy (anonymous):

use absolute after that limit

OpenStudy (anonymous):

can u explain a little further

OpenStudy (anonymous):

let f(x)= a(n)= (-1)^n (n+4/n+1) |f(x)|= lim |(-1)^n (n+4/n+1)| --> lim 1^n(n+4/n+1) = 1 \[\neq\] 0, means divergent

OpenStudy (anonymous):

Okay, alternating series test:\[\sum_{n=1}^{\infty}(-1)^n*a_n = a_1-a_2+a_3-a_4+...\] converges if limit of a_n as n approaches infinity is 0. The limit is 1 as n approaches infinity in (n+4)/(n+1), which does not equal 0, hence, the series diverges.

OpenStudy (anonymous):

thanks ya'll